OpenStudy (anonymous):

The sum of two numbers is 50 and the difference is 18 . What are the numbers?

OpenStudy (anonymous):

@NietzscheAlex

OpenStudy (jdoe0001):

numbers "a" and "b" their sum is \(\bf a + b = 50\) their difference is \(\bf a - b = 18\) solve by substitution

OpenStudy (jdoe0001):

if you were to solve for "b" in say a + b = 50 what would you end up with?